OpenStudy (anonymous):

can somebody help me with this system of equations. y+3=X 3X+4Y=16

OpenStudy (anonymous):

Plug the X from the top into the 2nd X \[3(y+3)+4y=16\]\[3y+9+4y=16\] Add common Units \[7y+9=16\]\[7y+9-9=16-9\] \[7y=7\]\[y=1\] Now put 1 in place of y \[1+3=x\]\[4=x\] \[(4,1)\]

OpenStudy (anonymous):

Is that what you were looking for?